### v1

@@ 2.4 Job Postings and Recruitment
An underlying assumption in our analysis is that job postings are an important source through which employers recruit workers and where workers learn about available jobs on the market. In order to assess the importance of job postings for recruitment, we link each ad in our vacancy data to a potential hire observed in our employer-employee data. Building this linkage requires us to assign a potential hire to each job ad by finding a job spell that starts (i) in the same establishment as the job posting, (ii) soon after the ad is posted, and (iii) in the same or a close occupation as the one assigned to the vacancy. We are able to link 89.4 percent of all publicly posted non-staffing agency job ads to a hire within the following six months; this applies to 94 percent for ads posted by public sector state employers, 90.2 percent of ads by non-state public sector employers (e.g., local municipalities, public schools, public hospitals), and 87.2 percent of ads by private sector employers. Conversely, 56.6 percent of all hires can be related to a posted ad within the previous six months; this applies to 90.1 percent of hires by public sector state employers, 62.8 percent of hires by non-state public employers, and 48 percent of private sector hires. [^18] While all employers have an obligation to report the vacancies they post to the Norwegian employment agency, it is up to them to decide whether or not to post a vacancy for the purposes of recruitment. For instance, employers may also use other recruitment channels, such as social networks, or directly solicit applications. However, recruitment by public sector state employers is regulated by state legislation, which requires as a general rule that all job openings are publicly advertised (the State Employees Act §4). This evidence suggests that job postings are indeed an important channel for job search and recruitment in the Norwegian labor market.
@@ 3 Pay and Non-Pay Attributes in Vacancy Texts
This section describes how we retrieve the job attributes publicly advertised by employers in job ads. These attributes are extracted from vacancy texts using tools developed in natural language processing. To the best of our knowledge, we are among the first to systematically extract a comprehensive set of job attributes from the texts of job ads.
@@ 3.1 Extracting Advertised Job Attributes
We identify commonly mentioned attributes in publicly posted vacancies using two distinct strategies: human recognition on a subset of job ads in our data set and identifying commonly used phrases in the “we offer” sections of vacancy texts. We first had several research assistants read through the text of 1,200 randomly chosen job ads in the vacancy corpus. The research assistants were asked to make a list of the commonly advertised attributes that are valuable to workers, to take note of the corresponding expressions associated with these attributes, and to indicate the presence of these attributes in each job posting.
As a second source of information on commonly advertised attributes, we isolate and extract common phrases from the “we offer” sections of job ads. We do this in two steps: First, we identify and extract all structured lists from the vacancy corpus. [^19] These lists are found by searching for consecutive sentences starting with a hyphen or bullet point and by searching for the HTML tags used to generate lists in online vacancies. By design, this approach identifies attributes that can be organised in structured lists. As an example, in the sample vacancy text in Figure 1, we retrieve three distinct lists: “Tasks,” “Required qualifications,” and “We can offer.” We find that 50 percent of the vacancy texts in our sample feature one or more such lists. Second, we isolate “we offer” lists by applying unsupervised topic modeling (Blei et al., 2003) to the collection of lists retrieved in the previous step. Further methodological details are provided in Appendix B. The most common expressions in the “we offer” lists constitute the second source of information on job attributes that we are able to identify.
Based on the two approaches described above, we are able to identify 47 distinct job attributes, as well as an initial set of expressions associated with each attribute. We summarize these 47 attributes in ten broad categories within classes of pay and non-pay job attributes, as shown in Table 2. Taken together, these attributes cover a variety of different job characteristics advertised in the text of job ads, such as information on career opportunities, convenient/inconvenient hours, and the quality of the workplace. The full list of job attributes with descriptions is provided in Appendix Tables B.3-B.5.
@@ A. Pay Attributes
   – Compensation Scheme Salary band; Bonuses; CBA-based pay
   – Financial Attributes Insurance scheme; Pension scheme
@@ B. Non-Pay Attributes
   – Hours of Work Full-time; Part-time; On-call job
   – Convenient Hours Regular daytime work; Flexible working hours
   – Inconvenient Hours Shift-work; Weekend/evening/nights
@@
   – Workplace Attributes Social environment; Remote work
@@
   – Other Minor Perks Location; Company vehicle; Company gym
Note: This table documents broad categories of attributes detected in vacancies, and examples of individual attributes. The full list of job attributes with descriptions is provided in Appendix Tables B.3-B.5.
@@
To expand on the set of possible expressions in our analysis, we rely on the Continuous Bag-of-Words (CBOW) Word2Vec algorithm (Mikolov et al., 2013). [^21] See Atalay et al. (2020) for an application of this algorithm to vacancy job titles. This approach measures the degree of similarity between phrases by exploiting the fact that words with similar meanings tend to occur in similar contexts. As an example, information about flexible working hours is commonly given with information about whether the spell is full-time or part-time in our corpus. We train the CBOW model using our collection of job ads, thus associating words that are similar in the context of job ads but not necessarily in other contexts. We then loop through the full collection of job ads and use the trained model to store phrases of one, two, or three consecutive words most similar to the phrases in our original lists from Step 1. Finally, we remove expressions that are wrongly classified as similar by the model. An example of a removed phrase is “flexible job adaptation,” which is classified as similar to “flexible working hours” by the model, but does not imply flexibility in one’s schedule. [^22] Notably, as there are two variants of written Norwegian (Bokmål and Nynorsk), we translate all phrases from Bokmål to Nynorsk, and append phrases that are not already captured by our dictionary. This discussion highlights the main advantage of our approach. We are able to identify a large number of attributes that can be described by many different expressions, but we still retain control on which specific expressions are included.
In total, our dictionary consists of 1,772 unique expressions, and each of the 47 attributes is associated with approximately 39 expressions on average, although some attributes have many fewer associated expressions than others. We have made the full list of expressions is available in an online repository, which is accessible through this link. Differences in the number of associated phrases reflect both the number of ways an attribute is presented in vacancy texts and the specificity of an attribute. For instance, the attribute “inclusive work-life scheme” is usually discussed using specific terms in Norwegian, which is why the model identifies only a small number of expressions for this attribute. Conversely, the attribute “shift work” is associated with many alternative expressions since there are many ways jobs can involve shift work.
@@
We simply use our 1,772 unique expressions to generate job attribute indicators for all ads in the data. A job ad is defined as advertising a given attribute if any of the expressions associated with that attribute is found in the job description.
@@ 3.2 Validation
A. Pay Attributes 67.1 79.2 73.2 89.0 89.3 96.6
   –– Compensation Scheme 53.1 64.8 61.8 92.5 91.9 96.4
   –– Financial Attributes 30.4 37.8 34.2 96.5 90.7 100.0
   –– Career Opportunities 41.1 48.8 41.8 79.0 71.3 83.2
B. Non-Pay Attributes 95.8 98.0 97.0 97.5 98.2 99.2
   –– Hours of Work 76.9 77.2 72.2 92.0 91.6 97.9
   –– Convenient Hours 20.3 15.8 15.8 94.0 81.0 81.0
   –– Inconvenient Hours 24.3 26.5 21.5 93.5 78.3 96.5
   –– Contract Duration 63.6 67.2 66.8 88.0 90.7 91.4
   –– Workplace Attributes 47.7 55.0 51.0 82.0 80.0 86.3
   –– Task-Related Attributes 52.6 65.0 46.5 72.0 64.2 89.8
   –– Minor Perks 39.5 46.0 18.5 66.5 33.7 83.8
Any Observed Attribute 96.4 98.2 98.0 98.2 99.0 99.2
Average Number of Attributes 5.84 6.64 5.56
Number of Job Ads 915,790 400 400 400
@@
Note: This table compares prevalence rates for ten broad categories of job attributes measured using our text analysis approach and manual recognition. A broad category of job attribute is considered present if at least one of the underlying distinct job attributes is detected. Column (1) is for the full sample of job ads posted between 2015 and 2019, while Columns (2)-(6) consider a random sample of 400 job ads used in the validation. Columns (1)-(3) report prevalence rates, while Columns (4)-(6) provide summary statistics that compare prevalence rates from text analysis and manual recognition. Success rate is defined as the share of job ads where the text analysis yields the same result as manual recognition. Precision is the rate of agreement between the two methods given that an attribute was detected in our text analysis, while, sensitivity is the rate of agreement given that an attribute was detected in the manual recognition. Precision is a measure of how many detected attributes are false positives, and sensitivity is a measure of how good our method is at recovering true attributes. See Appendix Table A.4 for results from validation exercises for each of the 47 underlying job attributes that contribute to the ten broad categories shown here.
@@
To check how important this limitation is in our setting, we performed an additional round of manual recognition. We selected another random sample of 400 job ads posted during the last five-year period, i.e., 2015-2019, and asked another group of research assistants to manually classify job attributes in this sample. We deliberately recruited a distinct group of assistants from the ones who helped us at the initial manual classification step (Step 1 of Section 3.1). We gave each research assistant our full list of 47 attributes and a few sample expressions for each attribute, and they were tasked with searching for these attributes in the full text of each job ad. We can then systematically compare the attributes retrieved with our automated procedure to those detected in the manual recognition performed by the assistants.
@@
Table 3 shows the results of this comparison for ten broad categories of job attributes. Column (1) reports the prevalence rates measured based on our text analysis for each of these categories in the full sample of job ads from 2015–2019 used in our analysis. Column (2) reports the corresponding rates for the random sample of 400 job ads, also based on our text analysis procedure, while Column (3) reports the prevalence rates from manual recognition for the random sample. Finally, Columns (4)-(6) provide rates of success, precision, and sensitivity by comparing prevalence rates for each category across text analysis and manual recognition.
@@ 3.3 The Prevalence of Advertised Job Attributes
We now use the detailed advertised job attributes to highlight several salient descriptive statistics. Figure 2 reports the share of job ads posted between 2015 and 2019 advertising each of the 47 attributes. About 55 percent of job ads provide some compensation-related information. However, explicit information about the actual level of pay is scarce. Less that 10 percent of ads mention a salary number or bracket (“compensation level”) and slightly more than 25 percent of ads indicate that pay is set by a collective wage bargaining agreement (“pay based on collective agreement”). [^24] By comparison, Batra et al. (2023) find that 13 percent of job ads have salary information in US data. Because these two attributes overlap to a large extent, 27 percent of job ads have information on the actual pay level of the position.
@@
Note: This figure documents the prevalence of job attributes detected in job ads posted in Norway between 2015 and 2019 [N=863,478]. The light blue bars show the share of ads detected having each of the distinct job attributes. The dark blue bars show the share of ads detected with at least one attribute within ten broad categories. Appendix Figure A.3 shows the corresponding results for the 2002-2019 period as a whole.
@@
Note: This figure shows pseudo $R^{2}$ (McFadden, 1974) from separate logistic regressions of binary job attributes detected in job ads posted in Norway between 2015 and 2019 [N=863,478] on fixed effects denoting unique combinations of 2-digit industries (90 groups), location indicators (10 groups), 2-digit occupations (43 groups), and 108,730 establishments. We start by including industry fixed effects, continue by including industry×$\times$×location fixed effects, and so on. The last set of regressions controls for 563,170 unique combinations. Location indicators group 422 municipalities into ten groups based on the number of workers such that municipalities assigned to the same group have a similar number of workers, with a specific indicator for job postings in Oslo. Appendix Figure A.4 shows the corresponding results for the 2002-2019 period as a whole.
@@
Next, we analyze to what extent differences in advertised attributes are explained by additional observable characteristics of the job ad, such as the industry of the posting establishment or the occupation associated with the job title stated in the job ad. In order to describe the variation in advertised job attributes that can be explained by such observable characteristics, we estimate a series of logistic regressions for the probability that a given attribute is advertised in the text of a vacancy with several sets of fixed effects: industries, locations, occupations, and establishments. Figure 3 reports the estimated pseudo-$R^{2}$, a summary measure of goodness of fit for binary outcomes (McFadden, 1974), from logistic regressions estimated separately for each attribute. [^25] As an alternative, we also estimated linear probability models using OLS with the same set of fixed effects separately for each attribute (results available upon request). However, as all of our pay and non-job attributes are binary, one could be concerned about interpreting standard goodness of fit from such regressions, which put restrictions on the range of the coefficient of determination $R^{2}$ (see, e.g., (Cox and Wermuth, 1992)). This measure therefore captures the variation in the content of job ads within industries, locations, occupations, and establishments.
A possibility is that the attributes advertised for a given position are largely determined by its industry or occupation, so the pay and non-pay attributes mentioned in vacancy texts directly reflect the type of job advertised. For example, it can be expected that working as a hospital nurse involves shift-work, in which case there is no value-added to mentioning this information in the corresponding job description. The data instead suggest that there is variation in advertised job attributes for very similar positions. Figure 3 shows that industries, occupations, and location fixed-effects explain at most 30 to 40 percent of the variation in the attributes advertised by employers across the 47 categories we retrieve. By contrast, when we add establishment fixed-effects, we find that the explained variation jumps up to 60 to 80 percent across attributes, which suggests that advertised attributes are in large part correlated across vacancies for establishments observed with several job ads. This last result is consistent with the existence of an “establishment fixed-effect” in advertised attributes, and it justifies the establishment-level analysis we develop in the next sections.

### v2

@@ 2.4 Job Postings and Recruitment
An underlying assumption in our analysis is that job postings are an important source through which employers recruit workers and where workers learn about available jobs. In order to assess the importance of job postings for recruitment, we link each ad in our vacancy data to a potential hire observed in our employer-employee data. Building this linkage requires us to assign a potential hire to each job ad by finding a job spell that starts (i) in the same establishment as the job posting, (ii) soon after the ad is posted, and (iii) in the same or a close occupation as the one in the ad. We are able to link 87.7 percent of all publicly posted non-staffing agency job ads to a hire within the following six months; this applies to 94.1 percent for ads posted by public sector state employers, 90.3 percent of ads by non-state public sector employers (e.g., local municipalities, public schools, public hospitals), and 87.5 percent of ads by private sector employers. Conversely, 42.6 percent of all hires can be related to a posted ad within the previous six months; this applies to 88.4 percent of hires by public sector state employers, 55.3 percent of hires by non-state public employers, and 35.7 percent of private sector hires. [^16] While all employers have an obligation to report the vacancies they post to the Norwegian employment agency, it is up to them to decide whether or not to post a vacancy for the purposes of recruitment. For instance, employers may also use other recruitment channels, such as social networks, or directly solicit applications. However, recruitment by public sector state employers is regulated by state legislation, which requires as a general rule that all job openings are publicly advertised (the State Employees Act §4). These numbers suggest that job postings are indeed an important channel for job search and recruitment in the Norwegian labor market.
@@ 3 Pay and Non-Pay Attributes in Vacancy Texts
This section describes how we retrieve the job attributes publicly advertised by employers in job ads. These attributes are extracted from vacancy texts using tools developed in natural language processing. To the best of our knowledge, we are among the first to systematically extract a comprehensive set of pay and amenity attributes from the texts of job ads.
@@ 3.1 Extracting Advertised Job Attributes
We identify commonly mentioned attributes in publicly posted vacancies using two distinct strategies: human recognition on a subset of job ads in our data set and identifying commonly used phrases in the “we offer” sections of vacancy texts. We first directed several research assistants read through the text of 1,200 randomly chosen job ads in the vacancy corpus. The research assistants were asked to make a list of the commonly advertised attributes that are valuable to workers, to take note of the corresponding expressions associated with these attributes, and to indicate the presence of these attributes in each job posting.
As a second source of information on commonly advertised attributes, we isolate and extract common phrases from the “we offer” sections of job ads. We do this in two steps: First, we identify and extract all structured lists from the vacancy corpus. [^17] These lists are found by searching for consecutive sentences starting with a hyphen or bullet point and by searching for the HTML tags used to generate lists in online vacancies. By design, this approach identifies attributes that can be organised in structured lists. In the later steps, we search for the common phrases identified in structured lists across the full corpus of job ads, including ads that were not posted online or lacked structured lists, as well as search for alternative similar phrases identified using text analysis tools. As an example, in the sample vacancy text in Figure 1, we retrieve three distinct lists: “Tasks,” “Required qualifications,” and “We can offer.” We find that 50 percent of the vacancy texts in our sample feature one or more such lists. Second, we isolate the “we offer” lists by applying unsupervised topic modeling (Blei et al., 2003) to the collection of lists retrieved in the previous step. Further methodological details are provided in Appendix B. The most common expressions in the “we offer” lists constitute the second source of information on commonly advertised job attributes that we are able to identify.
Based on the two approaches described above, we are able to identify 47 distinct job attributes, as well as an initial set of expressions associated with each attribute. For the purpose of illustration, we summarize the 47 attributes in ten broad categories within classes of pay and non-pay job attributes, as shown in Table 2. Taken together, these attributes cover a variety of different job characteristics advertised in job ads, such as information on career opportunities, convenient/inconvenient hours, and the quality of the workplace. The full list of job attributes with descriptions is provided in Appendix Tables B.3-B.5.
@@ A. Pay Attributes
   – Compensation Scheme Compensation level; Collective agreement pay; Incentive pay
   – Financial Attributes Insurance scheme; Pension scheme; Mortgage possibility
@@ B. Non-Pay Attributes
   – Hours of Work Full-time; Part-time; Full-time/part-time choice
   – Convenient Hours Regular daytime work; Possibility to work flexible hours
   – Inconvenient Hours Shift-work; Weekend/evening/nights; On-call employment
@@
   – Workplace Attributes Social environment; Good colleagues; Remote work
@@
   – Other Minor Perks Central location; Company vehicle; Company gym
Note: This table documents broad categories of attributes detected in vacancies, with notable examples of individual attributes. The full list of job attributes with descriptions is provided in Appendix Tables B.3-B.5.
@@
To expand on the set of possible expressions in our analysis, we rely on the Continuous Bag-of-Words (CBOW) Word2Vec algorithm (Mikolov et al., 2013). [^19] See Atalay et al. (2020) for an application of this algorithm to vacancy job titles. This approach measures the degree of similarity between phrases by exploiting that words with similar meanings tend to occur in similar contexts. As an example, information about flexible working hours is commonly given with information about whether the spell is full-time or part-time in our corpus. We train the CBOW model using our collection of job ads, thus associating words that are similar in the context of job ads but not necessarily in other contexts. We then loop through the full collection of job ads and use the trained model to store phrases of one, two, or three consecutive words most similar to the phrases in our original lists from Step 1. Finally, we remove expressions that are wrongly classified as similar by the model. An example of a removed phrase is “flexible job adaptation,” which is classified as similar to “flexible working hours” by the model, but does not imply flexibility in one’s schedule. [^20] Notably, as there are two variants of written Norwegian (Bokmål and Nynorsk), we translate all phrases from Bokmål to Nynorsk, and append phrases that are not already captured by our dictionary. This discussion highlights the main advantage of our approach. We are able to identify a large number of attributes that can be described by many different expressions, but we still retain control on which specific expressions are included.
In total, our dictionary consists of 1,772 unique expressions, and each of the 47 attributes is associated with around 39 expressions on average, although some attributes have many fewer associated expressions than others. We have made the full list of expressions is available in an online repository, which is accessible through this link. Differences in the number of associated phrases reflect both the number of ways an attribute is presented in vacancy texts and the specificity of an attribute. For instance, the attribute “inclusive work-life scheme” is usually discussed using specific terms in Norwegian, which is why the model identifies only a small number of expressions for this attribute. Conversely, the attribute “shift work” is associated with many alternative expressions since there are many ways jobs can involve shift work.
@@
We simply use our 1,772 unique expressions to generate job attribute indicators for all ads in the data. A job ad is defined as advertising a given attribute if any of the expressions associated with the attribute is found in the ad text.
@@ 3.2 Validation
A. Pay Attributes 70.8 79.2 73.2 89.0 89.3 96.6
   –– Compensation Scheme 55.9 64.8 61.8 92.5 91.9 96.4
   –– Financial Attributes 32.4 37.8 34.2 96.5 90.7 100.0
   –– Career Opportunities 43.6 48.8 41.8 79.0 71.3 83.2
B. Non-Pay Attributes 96.5 98.0 97.0 97.5 98.2 99.2
   –– Hours of Work 77.0 76.0 71.8 92.2 92.1 97.6
   –– Convenient Hours 20.5 15.8 15.8 94.0 81.0 81.0
   –– Inconvenient Hours 27.0 28.0 22.8 92.8 77.7 95.6
   –– Contract Duration 65.0 67.2 66.8 88.0 90.7 91.4
   –– Workplace Attributes 50.7 55.0 51.0 82.0 80.0 86.3
   –– Task-Related Attributes 55.9 65.0 46.5 72.0 64.2 89.8
   –– Minor Perks 39.8 46.0 18.5 66.5 33.7 83.8
Any Observed Attribute 97.0 98.2 98.0 98.2 99.0 99.2
Average Number of Attributes 6.11 6.64 5.56
Number of Job Ads 860,467 400 400 400
@@
Note: This table compares prevalence rates for ten broad categories of job attributes measured using our text analysis approach and manual recognition. A broad category of job attribute is considered present if at least one of the underlying distinct job attributes is detected. Column (1) is for the full sample of job ads with non-missing vacancy text posted between 2015 and 2019, while Columns (2)-(6) consider a random sample of 400 job ads used in the validation. Columns (1)-(3) report prevalence rates, while Columns (4)-(6) provide summary statistics that compare prevalence rates from text analysis and manual recognition. Success rate is defined as the share of job ads where the text analysis yields the same result as manual recognition. Precision is the rate of agreement between the two methods given that an attribute was detected in our text analysis, while, sensitivity is the rate of agreement given that an attribute was detected in the manual recognition. Precision is a measure of how many detected attributes are false positives, and sensitivity is a measure of how good our method is at recovering true attributes. See Appendix Table A.3 for results from validation exercises for each of the 47 underlying job attributes that contribute to the ten broad categories shown here.
@@
To check how important this concern is in our setting, we performed an additional round of manual recognition. We selected another random sample of 400 job ads posted during the last five-year period, i.e., 2015-2019, and directed another group of research assistants to manually classify job attributes in this sample. We deliberately recruited a distinct group of assistants from the ones who contributed to the initial manual classification (Step 1 of Section 3.1). We gave each research assistant our full list of 47 attributes and a few sample expressions for each attribute, and tasked them to search for each of these attributes in the full text of each job ad. We can then systematically compare the attributes retrieved with our automated procedure to those detected in the manual human recognition.
@@
Table 3 shows the results of this comparison for ten broad categories of job attributes. Column (1) reports the prevalence rates measured based on our text analysis for each of these categories in the sample of job ads from 2015–2019. Column (2) reports the corresponding rates for the random sample of 400 job ads, also based on our text analysis procedure, while Column (3) reports the prevalence rates from manual recognition for this random sample. Finally, Columns (4)-(6) provide rates of success, precision, and sensitivity by comparing prevalence rates for each category across text analysis and manual recognition.
@@ 3.3 The Prevalence of Advertised Job Attributes
We now use the detailed advertised job attributes to highlight several salient descriptive statistics. Figure 2 reports the share of job ads posted between 2015 and 2019 advertising each of the 47 attributes. About 55 percent of job ads provide some compensation-related information. However, explicit information about the actual level of pay is scarce. Less that 10 percent of ads mention a salary number or bracket (“compensation level”) and slightly more than 25 percent of ads indicate that pay is set by a collective wage bargaining agreement (“collective agreement pay”). [^22] By comparison, Batra et al. (2023) find that 13 percent of job ads have salary information in US data. As these two attributes can overlap, taken together, we find that 27 percent of job ads feature explicit information about the actual pay level.
@@
Note: This figure documents the prevalence of job attributes detected in job ads posted in Norway between 2015 and 2019 [N=860,467]. The light blue bars show the share of ads detected having each of the distinct job attributes. The dark blue bars show the share of ads detected with at least one attribute within ten broad categories. Appendix Figure A.3 shows the corresponding results for the 2002-2019 period as a whole.
@@
Note: This figure shows pseudo $R^{2}$ (McFadden, 1974) from separate logistic regressions of binary job attributes detected in job ads posted in Norway between 2015 and 2019 [N=719,538] on fixed effects denoting unique combinations of 2-digit industries (88 groups), location indicators (10 groups), 2-digit occupations (42 groups), and 59,259 establishments. We start by including industry fixed effects, continue by including industry×$\times$×location fixed effects, and so on. The last set of regressions controls for 114,380 unique combinations. The regression excludes 140,929 ads with a unique combination of establishment, occupation, location, and industry indicators, since these observations are perfectly predicted by the last set of regressions by construction. Location indicators group 422 municipalities into ten groups based on the number of workers such that municipalities assigned to the same group have a similar number of workers, with a specific indicator for job postings in Oslo. Appendix Figure A.4 shows the corresponding results for the 2002-2019 period as a whole.
@@
Next, we analyze to what extent differences in advertised attributes are explained by additional observable characteristics of the job ad, such as the industry of the posting establishment or the occupation associated with the job title stated in the job ad. To describe the variation in advertised job attributes that can be explained by these characteristics, we estimate a series of logistic regressions for the probability that a given attribute is advertised in the text of a vacancy with several sets of fixed effects: industries, locations, occupations, and establishments. Figure 3 reports the estimated pseudo-$R^{2}$, a summary measure of goodness of fit for binary outcomes (McFadden, 1974), from logistic regressions estimated separately for each attribute. [^23] As an alternative, we also estimated linear probability models using OLS with the same set of fixed effects separately for each attribute (results available upon request). However, as all of our pay and non-job attributes are binary, one could be concerned about interpreting standard goodness of fit from such regressions, which restrict the range of the coefficient of determination $R^{2}$ (Cox and Wermuth, 1992). This measure therefore captures the variation in the content of job ads within industries, locations, occupations, and establishments.
A possibility is that the attributes advertised for a given position are largely determined by its industry or occupation, so the pay and non-pay attributes mentioned in vacancy texts directly reflect the type of job advertised. For example, it can be expected that working as a hospital nurse involves shift-work, in which case there is no value-added to mentioning this information in the corresponding job description. The data instead suggest that there is variation in advertised job attributes for very similar positions. Figure 3 shows that industries, occupations, and location fixed-effects explain at most 30 to 40 percent of the variation in the attributes advertised by employers across the 47 categories we retrieve. By contrast, when we add establishment fixed-effects, we find that the explained variation jumps up to 50 to 70 percent across attributes, which suggests that advertised attributes are in large part correlated across vacancies for establishments observed with several job ads. This last result is consistent with the existence of an “establishment fixed-effect” in advertised attributes, and it justifies the establishment-level analysis we develop in the next sections.